# State-space models are accurate and efficient neural operators for dynamical systems

### v1

@@ A.1.3 Laplace Neural Operator (LNO)
Model 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
GRU 2.788E-2 1.033E-2 4.824E-3 2.418E-3 1.496E-3 8.716E-4 5.071E-4 2.283E-3 8.371E-4 2.716E-3
LSTM 2.204E-2 1.038E-2 5.291E-3 2.656E-3 2.640E-3 1.258E-3 5.251E-4 7.964E-4 2.297E-3 1.236E-3
DeepONet 3.078E-1 1.514E-1 7.831E-2 4.057E-2 2.216E-2 1.409E-2 1.047E-2 9.043E-3 8.742E-3 8.144E-3
FNO 2.069E-1 6.919E-2 2.324E-2 8.974E-3 3.875E-3 1.820E-3 9.467E-4 6.482E-4 4.485E-4 2.938E-4
LNO 2.016E-1 7.510E-2 2.667E-2 9.995E-3 4.069E-3 1.890E-3 9.395E-4 5.307E-4 3.206E-4 2.238E-4
Transformer 1.634E-1 6.393E-2 2.295E-2 8.857E-3 3.635E-3 1.697E-3 8.615E-4 5.278E-4 3.165E-4 2.073E-4
Oformer-V 2.481E-1 1.062E-1 4.016E-2 1.539E-2 5.934E-3 2.438E-3 1.130E-3 1.288E-3 1.244E-3 3.977E-4
Oformer-G 2.485E-1 1.102E-1 4.838E-2 2.112E-2 8.750E-3 4.292E-3 1.859E-3 1.522E-3 3.927E-3 3.120E-3
Oformer-F 1.191E+1 3.501E+0 1.406E+0 5.944E-1 2.482E-1 1.235E-1 7.293E-2 5.307E-2 4.145E-2 4.169E-2
GNOT 2.188E-2 1.129E-2 7.044E-3 3.848E-3 2.582E-3 1.723E-3 9.007E-4 1.353E-3 2.992E-3 2.000E-3
Mamba 3.161E-2 7.322E-3 2.743E-3 1.262E-3 6.692E-4 3.779E-4 2.502E-4 2.128E-4 1.547E-4 1.236E-4

### v2

@@ A.1.3 Laplace Neural Operator (LNO)
Model Metric 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
GRU Mean 2.751E-02 1.182E-02 5.528E-03 2.718E-03 1.775E-03 9.128E-04 5.157E-04 1.317E-03 1.242E-03 1.379E-03
Std 1.425E-02 6.274E-03 2.415E-03 1.005E-03 4.525E-04 1.638E-04 9.522E-05 6.882E-04 3.052E-04 9.488E-04
LSTM Mean 2.919E-02 1.221E-02 5.766E-03 2.948E-03 2.014E-03 1.111E-03 5.115E-04 8.236E-04 1.919E-03 1.023E-03
Std 5.071E-03 1.300E-03 3.408E-04 2.638E-04 4.978E-04 1.839E-04 1.073E-05 9.685E-05 7.694E-04 5.447E-04
DeepONet Mean 2.286E-01 1.068E-01 4.983E-02 2.191E-02 8.988E-03 4.229E-03 1.754E-03 1.569E-03 3.319E-03 2.724E-03
Std 1.583E-02 5.847E-03 3.733E-03 1.864E-03 6.847E-04 1.432E-04 1.285E-04 2.862E-04 4.322E-04 2.817E-04
FNO Mean 8.377E+00 2.224E+00 8.190E-01 3.350E-01 1.474E-01 7.468E-02 4.267E-02 2.874E-02 2.088E-02 2.121E-02
Std 1.056E+00 5.543E-01 1.901E-01 5.526E-02 1.419E-02 5.552E-03 5.784E-03 5.938E-03 6.394E-03 5.104E-03
LNO Mean 1.721E-02 8.301E-03 4.883E-03 2.734E-03 1.676E-03 1.485E-03 6.643E-04 9.606E-04 2.757E-03 2.011E-03
Std 4.271E-03 2.477E-03 1.695E-03 9.095E-04 6.982E-04 5.560E-04 2.163E-04 3.904E-04 1.260E-03 9.660E-04
Transformer Mean 6.392E-01 2.523E-01 1.119E-01 5.659E-02 3.253E-02 2.225E-02 1.714E-02 1.487E-02 1.349E-02 1.260E-02
Std 1.094E-02 3.294E-03 1.744E-03 2.322E-03 1.528E-03 1.184E-03 1.184E-03 1.219E-03 1.046E-03 1.021E-03
Oformer-V Mean 4.515E-01 1.485E-01 4.930E-02 1.877E-02 7.960E-03 3.741E-03 1.940E-03 1.141E-03 8.460E-04 5.243E-04
Std 5.774E-02 3.174E-02 1.342E-02 5.487E-03 2.479E-03 1.187E-03 6.441E-04 3.338E-04 2.280E-04 1.732E-04
Oformer-G Mean 2.647E-01 7.146E-02 2.095E-02 7.286E-03 2.848E-03 1.278E-03 6.462E-04 4.302E-04 2.877E-04 1.816E-04
Std 3.578E-02 4.720E-03 6.421E-04 6.479E-04 3.276E-04 1.686E-04 9.763E-05 6.990E-05 4.816E-05 2.007E-05
Oformer-F Mean 2.314E-01 6.783E-02 2.109E-02 7.714E-03 3.195E-03 1.490E-03 7.633E-04 5.409E-04 3.762E-04 2.156E-04
Std 3.078E-02 3.636E-03 1.555E-03 7.555E-04 3.062E-04 1.596E-04 9.401E-05 3.987E-05 1.575E-05 1.008E-05
GNOT Mean 1.552E-01 5.804E-02 2.163E-02 8.720E-03 3.855E-03 1.753E-03 8.726E-04 6.241E-04 3.887E-04 2.482E-04
Std 4.856E-03 3.312E-03 1.625E-03 9.031E-04 6.404E-04 2.752E-04 1.442E-04 4.299E-05 5.183E-05 2.275E-05
Mamba Mean 3.091E-02 8.101E-03 3.108E-03 1.450E-03 7.212E-04 3.993E-04 2.571E-04 2.233E-04 1.733E-04 1.352E-04
Std 2.409E-03 2.090E-03 1.030E-03 4.888E-04 1.479E-04 7.192E-05 2.083E-05 1.281E-05 1.876E-05 9.883E-06